Colt Company owns a machine that can produce two specialized products. Production time for Product TLX is two units per hour and for Product MTV is four units per hour. The machine’s capacity is 2,500 hours per year. Both products are sold to a single customer who has agreed to buy all of the company’s output up to a maximum of 4,250 units of Product TLX and 1,910 units of Product MTV. Selling prices and variable costs per unit to produce the products follow. Product TLX Product MTV

Selling price per unit $12.50 $7.50 Variable costs per unit 3.75 4.50

Determine the company's most profitable sales mix and the contribution margin that results from that sales mix. (Round cost per unit answers to 2 decimal places.)
Answer #1

Contribution margin per hour

TLX MTV
Contribution margin per unit 8.75 3
Hour per unit 0.50 0.25
Contribution margin per hour 17.50 12

Calculate following

Hour Unit Contribution margin
TLX 4250*.50 = 2125 4250 2125*17.50 = 37187.50
MTV 2500-2125 = 375 375*4 = 1500 375*12 = 4500
TOtal 41687.50 or 41688